Two pipes are used to fill an oil tanker. When the larger pipe is used alone, it takes 60 hours to fill the tanker. When the smaller pipe is used alone, it takes 80 hours to fill the tanker. The large pipe begins filling the tanker. After 20 hours, the large pipe is closed and the smaller pipe is opened. How much longer will it take to finish filling the tanker using only the smaller pipe?
Answer by mananth(16949)

Two pipes are used to fill an oil tanker. When the larger pipe is used alone, it takes 60 hours to fill the tanker. When the smaller pipe is used alone, it takes 80 hours to fill the tanker. The large pipe begins filling the tanker. After 20 hours, the large pipe is closed and the smaller pipe is opened. How much longer will it take to finish filling the tanker using only the smaller pipe?
..
larger pipe 60 hours
smaller pipe 80 hours
..
Larger pipe does 1/60 of the job in 1 hour
so in 20 hours it does 20/60 of the job .= 1/3 of the job.
..
balance job will be 1-1/3 = 2/3 of the job.
..
smaller pipe does 1/80 of the job in 1 hour
for 2/3 of the job it will take ?
(2/3) / (1/80)
2/3 *80
=160/3
=53.33 hours to do the remaining job of filling the tank.
...
m.ananth@hotmail.ca